OpenStudy (anonymous):

In a game of blackjack, a 2-card hand consisting of an ace and a face card or a 10 is called a blackjack. (Round your answers to four decimal places.) (a) If a player is dealt 2 cards from a standard deck of 52 well-shuffled cards, what is the probability that the player will receive a blackjack? (b) If a player is dealt 2 cards from 2 well shuffled standard decks, what is the probability that the player will receive a blackjack?

OpenStudy (ybarrap):

4 aces 3 face cards per suit 1 10 card per suit 4 suits ways to get blackjack: 1) face card and an ace = (4choose1)*(3choose1)*(4choose1) 2) 10 card and an ace = (4choose1)*(4choose1) 3) total number of ways to deal a hand (52 choose 2) P(blackjack)= (P(1) + P(2))/P(3) = ((4choose1)^2(3choose1 + 1)/(52choose2) = 4^2(3 + 1)/(52*51/2) = 4^3/51/26 = .04827